OpenStudy (anonymous):

If f(x) = 3 sin x + ln(4x), find f '(x). f '(x) =

OpenStudy (anonymous):

f(x) = 3 sin x + ln(4x) f'(x) = d/dx (3 sin x) + d/dx (ln(4x)) d/dx sinx = cosx d/dx lnx = 1/x Can you do it now?

OpenStudy (anonymous):

3cosx+....

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

d/dx ln(ax) = (1/ax) (a) = ...

OpenStudy (anonymous):

3cosx+1/x right??

OpenStudy (anonymous):

It should be!

OpenStudy (anonymous):

I meant it is correct.

OpenStudy (anonymous):

thanks